OpenStudy (anonymous):

If two marbles are removed from the jar, what is the probability that it is red, white, or blue?

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

I forgot

OpenStudy (anonymous):

Yup, the rest :P

OpenStudy (anonymous):

a jar contains 5 marbles, 1 of each of the colors red, white, blue, green, and yellow

OpenStudy (anonymous):

we are referring to that

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

Well I have two possibilities, are they both taking out at the same time, or one afyer the other?

OpenStudy (anonymous):

Mhmm... it should still be the same thing anyways.

OpenStudy (anonymous):

3/10 is the answer

OpenStudy (chihiroasleaf):

I think, first you have to find the number of ways to take 2 marbles out of 5 marbles ( n(S)) then find number of ways to take 2 marbles (which is red, white or blue out of 5 marbles, it can be, red-white, white-blue, red-blue, ( n(A) ) P(A) = n(A) / n(S)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

dang I misread the question!

OpenStudy (chihiroasleaf):

you can use combination @luisz

OpenStudy (anonymous):

let us see what jiji has to say

OpenStudy (anonymous):

You can get 3 colors.... so: 3/5 Next, there is only 3 marbles left and 2 colors you can get. 2/4 3/5 x 2/4 = 6/20 or 3/10

OpenStudy (anonymous):

4 marbles left****

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

after he picked the first one, he picked another one

OpenStudy (anonymous):

the probability that it is red, white, or blue? Taht means just get any of the 3.

OpenStudy (anonymous):

3/5 times 2/4=6/20

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

ty

OpenStudy (chihiroasleaf):

I use this way, A = 3 (red-white, red-blue, white-blue) S = 5C3 = 10 so, the possibilities is 3/10

OpenStudy (anonymous):

That works also.